Accelerator Update for February 16, 2009 - February 18, 2009
Monday Feb 16
The day shift began with Operations monitoring the Tevatron (TeV) store 6810 (240.85E30) and the Recycler stash (104.6E10), with the Antiproton Source stacking (26.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
Operations disabled the SY beam at 9:52 AM.
At 10:54 AM, a Linac RF station (LRF3) tripped off and would not reset. Linac experts investigated.
A Pbar expert began a study at 2:04 PM that lowered its intensity.
A MI expert began a coalescing check at 2:51 PM.
TeV experts began separator scans at 3:07 PM.
Operations reported at 4:04 PM that the Recycler had stashed antiprotons seven times during the day shift.
At 7:33 PM, MIRF14 tripped off for the first of three times in about three hours. It reset.
Operations terminated store 6810 at 8:38 PM.
Operations put the TeV into shot setup at 8:40 PM.
At 8:56 PM, the P1 transfer line beam permit tripped off when the Pbar autotune program set a power supply (HF13) to a wrong output. This happened three times. Operations restored the program with a different save file.
At 9:16 PM, MIRF14 tripped off. It reset.
A TeV expert began a 20-minute tune-up on the tune an chromaticity tracker at 9:28 PM.
Operations established store 6811 at 10:53 PM with and initial luminosity of 283.0E30.
Operations reported at midnight that the Recycler had stashed six times during the evening shift.
Tuesday Feb 17
The midnight shift began with Operations monitoring store 6811 (217.74E30) and the stash (83.6E10), with Pbar stacking (16.0E10), and with NuMI and MiniBooNE taking beam.
At 2:21 AM, LRF3 drifted low. It reset.
At 3:47 AM, the Linac low energy buncher tripped off for the first of two times in about two hours. It reset.
Operations reported at 7:54 AM that the Recycler had stashed nine times during the midnight shift.
Operations prepared M01 for a one-hour access at 8:35 AM.
At 11:05 AM, Operations reported MI beam permit trips due to bad beam positions. Later, operators reported that a Booster power supply (SEXSB) was off and causing an emittance blow up. Booster experts turned the supply on and the trips stopped.
Operations terminated store 6811 at 2:37 PM.
Operations put the TeV into shot setup at 2:49 PM.
Operations reported at 4 PM that the Recycler had stashed five times during the day shift.
Operations established store 6812 at 5 PM with an initial luminosity of 307.1E30.
At 8:19 PM, Operations reported that the first transfer of a Recycler stash didn’t make it. Operators investigated and discovered a single instance of a timing problem. The rest of the transfers were normal.
A Pbar expert began the first of two beam measurements at 9:18 PM that required turning off the stacktail monitor.
Operations reported at midnight that the Recycler had stashed eight times during the evening shift.
Wednesday Feb 18
The midnight shift began with Operations monitoring store 6812 (119.84E30) and the stash (221.7E10), with Pbar stacking (15.9E10), and with NuMI and MiniBooNE taking beam.
At 12:37 AM, BRF16 tripped off for the first of four times in one minute. An operator adjusted the Anode Program.
At 2:53 AM, a Linac quadrupole power supply (L:QPS224) tripped off. It would not reset. Operations called in an expert.
A Linac expert reported at 5:12 AM that he had replaced the L:QPS224 power supply. Linac resumed sending beam.
At 7:03 AM, the Pelletron tripped off. It reset.
Operations reported at 8:04 AM that the Recycler had stashed six times during the midnight shift.
The Plans for Wednesday
The plans for today are to stack, stash, and store. The Run Coordinator announced that the Pbar group would be replacing their target on Thursday morning (2/19/09) beginning around 9 AM. The TeV will maintain a store and the Recycler its stash. Other machines may access their areas for maintenance.
Comments and Suggestions
|last modified 02/18/2009 email Fermilab|
|Security, Privacy, Legal|